OpenStudy (anonymous):

Explain the difference between using the cosine ratio to solve for a missing angle in a right triangle versus using the secant ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view.

OpenStudy (anonymous):

@mathstudent55 @mathmale

OpenStudy (anonymous):



OpenStudy (anonymous):

How do I write that in paragraph form though

OpenStudy (anonymous):

I need help with this last one too

OpenStudy (anonymous):

\[\tan 49=\frac{ h }{900 },h=900*\tan 49=?\]

OpenStudy (anonymous):

thats the last question I just sent?

OpenStudy (anonymous):

@surjithayer what would be the asnwer?

OpenStudy (anonymous):

h=900*1.15=1035 ft

OpenStudy (anonymous):

THANKS SO MUCH